# Improving Robustness to Model Inversion Attacks via Sparse Coding Architectures

### v1

@@ 0.A.3 Cluster Details
We run all our experiments using the slurm batch jobs on industry-standard high-performance GPU clusters with $40$ cores and $4$ nodes. Details of the hardware and architecture of our cluster are described in Table 4. We note that noise-based Gaussian and Titcombe et al. [71] defenses are typically fastest on this architecture (though they are the least-performant). We emphasize that our sparse coding implementations are ‘research-grade’, unlike the optimized torch GAN implementations available for [22]. See also Appendix 0.F. Note that for large scale applications, SCA’s sparse coding updates can be accelerated such that they can be computed extremely efficiently (see the training complexity discussion at the end of Section 3).
@@ 0.A.5 Attack details
We vary the sparsity, i.e., $\lambda$ parameter and run the Sparse-Standard, as well as our SCA. We observe that increasing $\lambda$ helps improve the robustness, without significant accuracy drops. For example, Table  0.A.6 shows this comparison for MNIST in the end-to-end setting.
@@ Appendix 0.B Model Inversion Attack Methodology: Additional discussion
Because privacy attacks are an emerging field, we feel it is relevant to include additional context and discussion here. Recent work has highlighted a variety of attack vectors targeting sensitive training data of machine learning models [46, 18, 76, 72, 66, 90, 15, 17, 75, 61, 21, 93, 12, 74, 44, 11]. Adversaries with different access (i.e., black-box, white-box) to these models perform different attacks leveraging a wide range of capabilities, e.g., knowledge about the target model confusion matrix and access to blurred images of that particular class [20, 15, 26, 77, 34]. Such attacks commonly fall under the umbrella of privacy attacks, which include specific attacker goals such as membership inference, model stealing, model inversion, etc. [49, 80, 88, 31].
@@ Appendix 0.C Results of extra {threat model, dataset} experiments
Accuracy
MNIST
No-Defense
$7.24$
$0.783$
$23.6$
$0.971$
Gaussian-Noise
$6.94$
$0.686$
$31.22$
$0.958$
GAN
$6.83$
$0.734$
$89.38$
$0.968$
Gong et al.[22]++
$6.69$
$0.716$
$92.21$
$0.987$
Titcombe et al.[71]
$6.34$
$0.744$
$131.8$
$0.980$
Gong et al.[22]
$6.76$
$0.681$
$99.53$
$0.985$
Peng et al.[58]
$6.89$
$0.704$
$283.8$
$0.941$
Hayes et al.[25]
$7.03$
$0.672$
$396.1$
$0.871$
Wang et al.[78]
$7.14$
$0.752$
$261.2$
$0.937$
Sparse-Standard
$6.24$
$0.631$
$158.6$
$0.986$
SCA0.1
$\mathbf{6.19}$
$\mathbf{0.633}$
$\mathbf{287.9}$
$\mathbf{0.984}$
SCA0.25
$\mathbf{5.83}$
$\mathbf{0.607}$
$\mathbf{289.3}$
$\mathbf{0.983}$
SCA0.5
$\mathbf{5.74}$
$\mathbf{0.604}$
$\mathbf{299.6}$
$\mathbf{0.977}$
Fashion
No-Defense
$8.91$
$0.147$
$235.5$
$0.886$
MNIST
Gaussian-Noise
$8.67$
$0.132$
$239.8$
$0.815$
GAN
$8.66$
$0.147$
$243.3$
$0.883$
Gong et al.[22]++
$8.73$
$0.130$
$220.2$
$0.906$
Titcombe et al.[71]
$8.56$
$0.134$
$229.8$
$0.905$
Gong et al.[22]
$8.57$
$0.143$
$244.3$
$0.888$
Peng et al.[58]
$8.85$
$0.147$
$227.5$
$0.845$
Hayes et al.[25]
$8.63$
$0.139$
$218.4$
$0.752$
Wang et al.[78]
$8.90$
$0.119$
$210.3$
$0.880$
Sparse-Standard
$8.71$
$0.135$
$223.3$
$0.879$
SCA0.1
$\mathbf{8.49}$
$\mathbf{0.039}$
$\mathbf{222.8}$
$\mathbf{0.897}$
SCA0.25
$\mathbf{8.49}$
$\mathbf{0.032}$
$\mathbf{229.9}$
$\mathbf{0.887}$
SCA0.5
$\mathbf{8.45}$
$\mathbf{0.047}$
$\mathbf{233.5}$
$\mathbf{0.876}$
CIFAR10
No-Defense
$11.94$
$0.381$
$39.38$
$0.821$
Gaussian-Noise
$11.88$
$0.365$
$77.92$
$0.626$
GAN
$11.86$
$0.369$
$88.39$
$0.596$
Titcombe et al.[71]
$10.89$
$0.346$
$79.19$
$0.792$
Gong et al.[22]++
$11.06$
$0.339$
$78.48$
$0.773$
Gong et al.[22]
$11.21$
$0.334$
$92.33$
$0.682$
Peng et al.[58]
$11.96$
$0.354$
$120.5$
$0.752$
Hayes et al.[25]
$11.12$
$0.342$
$142.1$
$0.626$
Wang et al.[78]
$11.02$
$0.346$
$142.6$
$0.756$
Sparse-Standard
$10.74$
$0.303$
$137.4$
$0.790$
SCA0.1
$\mathbf{10.59}$
$\mathbf{0.305}$
$\mathbf{144.1}$
$\mathbf{0.787}$
SCA0.25
$\mathbf{10.27}$
$\mathbf{0.279}$
$\mathbf{189.9}$
$\mathbf{0.772}$
SCA0.5
$\mathbf{10.23}$
$\mathbf{0.276}$
$\mathbf{189.7}$
$\mathbf{0.744}$
Dataset
Defense
@@
[…]
Accuracy
MNIST
No-Defense
$40.87$
$0.982$
$16.31$
$0.971$
Gaussian-Noise
$40.88$
$0.983$
$15.88$
$0.958$
GAN
$40.69$
$0.981$
$16.59$
$0.968$
Titcombe et al.[71]
$31.18$
$0.863$
$47.32$
$0.980$
Gong et al.[22]++
$30.37$
$0.838$
$72.99$
$0.987$
Gong et al.[22]
$29.05$
$0.817$
$75.39$
$0.985$
Peng et al.[58]
$18.44$
$0.354$
$111.6$
$0.968$
Hayes et al.[25]
$19.75$
$0.488$
$298.8$
$0.871$
Wang et al.[78]
$27.26$
$0.862$
$72.66$
$0.962$
Sparse-Standard
$21.34$
$0.439$
$142.9$
$0.986$
SCA0.1
$\mathbf{19.54}$
$\mathbf{0.502}$
$\mathbf{178.5}$
$\mathbf{0.984}$
SCA0.25
$\mathbf{18.81}$
$\mathbf{0.340}$
$\mathbf{174.1}$
$\mathbf{0.983}$
SCA0.5
$\mathbf{17.85}$
$\mathbf{0.164}$
$\mathbf{335.5}$
$\mathbf{0.977}$
Fashion
No-Defense
$37.86$
$0.975$
$13.91$
$0.886$
MNIST
Gaussian-Noise
$36.54$
$0.969$
$16.49$
$0.815$
GAN
$37.68$
$0.974$
$19.26$
$0.883$
Gong et al.[22]++
$27.71$
$0.794$
$41.35$
$0.906$
Titcombe et al.[71]
$26.66$
$0.759$
$53.76$
$0.905$
Gong et al.[22]
$21.24$
$0.523$
$93.08$
$0.888$
Peng et al.[58]
$17.98$
$0.368$
$70.53$
$0.880$
Hayes et al.[25]
$21.13$
$0.297$
$223.3$
$0.752$
Wang et al.[78]
$25.98$
$0.806$
$41.87$
$0.838$
Sparse-Standard
$19.35$
$0.446$
$128.4$
$0.879$
SCA0.1
$\mathbf{17.92}$
$\mathbf{0.209}$
$\mathbf{196.1}$
$\mathbf{0.897}$
SCA0.25
$\mathbf{17.03}$
$\mathbf{0.186}$
$\mathbf{195.2}$
$\mathbf{0.887}$
SCA0.5
$\mathbf{14.51}$
$\mathbf{0.069}$
$\mathbf{423.2}$
$\mathbf{0.876}$
CIFAR10
No-Defense
$21.17$
$0.477$
$70.96$
$0.821$
Gaussian-Noise
$20.26$
$0.220$
$77.42$
$0.626$
GAN
$19.71$
$0.259$
$132.0$
$0.596$
Titcombe et al.[71]
$18.62$
$0.174$
$171.9$
$0.792$
Gong et al.[22]++
$18.27$
$0.209$
$149.1$
$0.773$
Gong et al.[22]
$19.10$
$0.150$
$133.8$
$0.682$
Peng et al.[58]
$17.20$
$0.002$
$130.3$
$0.717$
Hayes et al.[25]
$17.95$
$0.002$
$142.4$
$0.626$
Wang et al.[78]
$17.08$
$0.002$
$136.1$
$0.793$
Sparse-Standard
$18.01$
$0.003$
$168.6$
$0.790$
SCA0.1
$\mathbf{17.09}$
$\mathbf{0.001}$
$\mathbf{172.0}$
$\mathbf{0.787}$
SCA0.25
$\mathbf{16.78}$
$\mathbf{0.001}$
$\mathbf{189.3}$
$\mathbf{0.772}$
SCA0.5
$\mathbf{16.24}$
$\mathbf{0.001}$
$\mathbf{197.0}$
$\mathbf{0.744}$
Dataset
Defense
[…]
Accuracy
MNIST
No-Defense
$31.21$
$0.923$
$19.64$
$0.963$
Gaussian-Noise
$31.07$
$0.922$
$23.27$
$0.972$
GAN
$28.39$
$0.894$
$27.26$
$0.969$
Gong et al.[22]
$28.30$
$0.806$
$69.38$
$0.986$
Titcombe et al.[71]
$25.40$
$0.713$
$76.88$
$0.952$
Gong et al.[22]++
$21.94$
$0.591$
$97.33$
$0.991$
Peng et al.[58]
$16.90$
$0.475$
$103.2$
$0.960$
Hayes et al.[25]
$17.23$
$0.030$
$288.1$
$0.856$
Wang et al.[78]
$21.87$
$0.696$
$53.09$
$0.903$
Sparse-Standard
$18.71$
$0.288$
$188.4$
$0.981$
SCA0.1
$\mathbf{16.17}$
$\mathbf{0.109}$
$\mathbf{227.4}$
$\mathbf{0.988}$
SCA0.25
$\mathbf{17.40}$
$\mathbf{0.058}$
$\mathbf{301.6}$
$\mathbf{0.980}$
SCA0.5
$\mathbf{14.98}$
$\mathbf{0.044}$
$\mathbf{307.7}$
$\mathbf{0.975}$
Fashion No-Defense
$29.66$
$0.911$
$14.33$
$0.868$
MNIST
Gaussian-Noise
$29.49$
$0.909$
$14.81$
$0.871$
GAN
$26.03$
$0.849$
$19.33$
$0.885$
Gong et al.[22]
$23.70$
$0.631$
$97.52$
$0.884$
Titcombe et al.[71]
$20.48$
$0.565$
$81.01$
$0.872$
Gong et al.[22]++
$25.77$
$0.726$
$57.72$
$0.908$
Peng et al.[58]
$20.67$
$0.583$
$46.48$
$0.865$
Hayes et al.[25]
$20.10$
$0.256$
$200.6$
$0.748$
Wang et al.[78]
$24.53$
$0.588$
$81.79$
$0.881$
Sparse-Standard
$19.54$
$0.405$
$200.5$
$0.882$
SCA0.1
$\mathbf{18.11}$
$\mathbf{0.154}$
$\mathbf{171.1}$
$\mathbf{0.904}$
SCA0.25
$\mathbf{17.74}$
$\mathbf{0.188}$
$\mathbf{203.8}$
$\mathbf{0.896}$
SCA0.5
$\mathbf{17.15}$
$\mathbf{0.134}$
$\mathbf{270.4}$
$\mathbf{0.879}$
CIFAR10
No-Defense
$16.48$
$0.709$
$47.77$
$0.823$
Gaussian-Noise
$14.79$
$0.311$
$149.5$
$0.598$
GAN
$14.87$
$0.296$
$13.01$
$0.675$
Titcombe et al.[71]
$14.68$
$0.244$
$157.3$
$0.779$
Gong et al.[22]++
$13.32$
$0.003$
$162.4$
$0.691$
Gong et al.[22]
$14.55$
$0.291$
$152.1$
$0.644$
Peng et al.[58]
$17.18$
$0.002$
$169.1$
$0.707$
Hayes et al.[25]
$15.44$
$0.005$
$204.5$
$0.596$
Wang et al.[78]
$14.73$
$0.001$
$176.3$
$0.820$
Sparse-Standard
$13.22$
$0.003$
$167.9$
$0.769$
SCA0.1
$\mathbf{13.18}$
$\mathbf{0.002}$
$\mathbf{174.2}$
$\mathbf{0.758}$
SCA0.25
$\mathbf{13.07}$
$\mathbf{0.002}$
$\mathbf{181.2}$
$\mathbf{0.742}$
SCA0.5
$\mathbf{12.88}$
$\mathbf{0.002}$
$\mathbf{375.3}$
$\mathbf{0.739}$
Dataset
Defense
[…]
Accuracy
MNIST
Titcombe et al.[71]-$1.0$
$6.60$
$0.685$
$280.1$
$0.938$
SCA0.1
$\mathbf{6.19}$
$\mathbf{0.633}$
$\mathbf{287.9}$
$\mathbf{0.984}$
SCA0.25
$\mathbf{5.83}$
$\mathbf{0.607}$
$\mathbf{289.3}$
$\mathbf{0.983}$
SCA0.5
$\mathbf{5.74}$
$\mathbf{0.604}$
$\mathbf{299.6}$
$\mathbf{0.977}$
Fashion
Titcombe et al.[71]-$1.0$
$8.72$
$0.1412$
$232.1$
$0.823$
MNIST
SCA0.1
$\mathbf{8.49}$
$\mathbf{0.039}$
$\mathbf{222.8}$
$\mathbf{0.897}$
SCA0.25
$\mathbf{8.49}$
$\mathbf{0.032}$
$\mathbf{229.9}$
$\mathbf{0.887}$
SCA0.5
$\mathbf{8.45}$
$\mathbf{0.047}$
$\mathbf{233.5}$
$\mathbf{0.876}$
CIAFR10
Titcombe et al.[71]-$1.0$
$10.75$
$0.335$
$112.7$
$0.779$
SCA0.1
$\mathbf{10.59}$
$\mathbf{0.305}$
$\mathbf{144.1}$
$\mathbf{0.787}$
SCA0.25
$\mathbf{10.27}$
$\mathbf{0.279}$
$\mathbf{189.9}$
$\mathbf{0.772}$
SCA0.5
$\mathbf{10.23}$
$\mathbf{0.276}$
$\mathbf{189.7}$
$\mathbf{0.744}$
Dataset
Defense
[…]
Accuracy
MNIST
Titcombe et al.[71]-$1.0$
$24.89$
$0.664$
$50.64$
$0.938$
SCA0.1
$\mathbf{19.54}$
$\mathbf{0.502}$
$\mathbf{178.5}$
$\mathbf{0.984}$
SCA0.25
$\mathbf{18.81}$
$\mathbf{0.340}$
$\mathbf{174.1}$
$\mathbf{0.983}$
SCA0.5
$\mathbf{17.85}$
$\mathbf{0.164}$
$\mathbf{335.5}$
$\mathbf{0.977}$
Fashion
Titcombe et al.[71]-$1.0$
$20.21$
$0.567$
$80.55$
$0.823$
MNIST
SCA0.1
$\mathbf{17.92}$
$\mathbf{0.209}$
$\mathbf{196.1}$
$\mathbf{0.897}$
SCA0.25
$\mathbf{17.03}$
$\mathbf{0.186}$
$\mathbf{195.2}$
$\mathbf{0.887}$
SCA0.5
$\mathbf{14.51}$
$\mathbf{0.069}$
$\mathbf{423.2}$
$\mathbf{0.876}$
CIFAR10
Titcombe et al.[71]-$1.0$
$18.71$
$0.672$
$170.8$
$0.779$
SCA0.1
$\mathbf{17.09}$
$\mathbf{0.001}$
$\mathbf{172.0}$
$\mathbf{0.787}$
SCA0.25
$\mathbf{16.78}$
$\mathbf{0.001}$
$\mathbf{189.3}$
$\mathbf{0.772}$
SCA0.5
$\mathbf{16.24}$
$\mathbf{0.001}$
$\mathbf{197.0}$
$\mathbf{0.744}$

### v2

@@ 0.A.3 Cluster Details
We run all our experiments using the slurm batch jobs on industry-standard high-performance GPU clusters with $40$ cores and $4$ nodes. Details of the hardware and architecture of our cluster are described in Table 4. We note that noise-based Gaussian and Titcombe et al. [75] defenses are typically fastest on this architecture (though they are the least-performant). We emphasize that our sparse coding implementations are ‘research-grade’, unlike the optimized torch GAN implementations available for [25]. See also Appendix 15. Note that for large scale applications, SCA’s sparse coding updates can be accelerated such that they can be computed extremely efficiently (see the training complexity discussion in the main paper body).
@@ 0.A.5 Attack details
We vary the sparsity, i.e., $\lambda$ parameter and run the Sparse-Standard, as well as our SCA. We observe that increasing $\lambda$ helps improve the robustness, without significant accuracy drops. For example, Table  6 shows this comparison for MNIST in the end-to-end setting.
@@ Appendix 0.B Model Inversion Attack Methodology: Additional discussion
Because privacy attacks are an emerging field, we feel it is relevant to include additional context and discussion here. Recent work has highlighted a variety of attack vectors targeting sensitive training data of machine learning models [49, 19, 80, 76, 70, 95, 16, 18, 19, 79, 65, 24, 98, 13, 78, 47, 12]. These attacks not only target centralized models but also can make the federated learning models vulnerable to attacks [29, 21]. Adversaries with different access (i.e., black-box, white-box) to these models perform different attacks leveraging a wide range of capabilities, e.g., knowledge about the target model confusion matrix and access to blurred images of that particular class [23, 16, 29, 81, 37]. Such attacks commonly fall under the umbrella of privacy attacks, which include specific attacker goals such as membership inference, model stealing, model inversion, etc. [52, 84, 93, 34]. Defending against privacy attacks is a core task of mainstream technology platforms ranging from public social networks to private medical research [9, 57, 90].
@@ Appendix 0.C Results of extra {threat model, dataset} experiments
MNIST No-Defense $7.24$ $0.783$ $23.6$ $0.971$
Gaussian-Noise $6.94$ $0.686$ $31.22$ $0.958$
GAN $6.83$ $0.734$ $89.38$ $0.968$
Gong et al.[25]++ $6.69$ $0.716$ $92.21$ $0.987$
Titcombe et al.[75] $6.34$ $0.744$ $131.8$ $0.980$
Gong et al.[25] $6.76$ $0.681$ $99.53$ $0.985$
Peng et al.[62] $6.89$ $0.704$ $283.8$ $0.941$
Hayes et al.[28] $7.03$ $0.672$ $396.1$ $0.871$
Wang et al.[82] $7.14$ $0.752$ $261.2$ $0.937$
Sparse-Standard $6.24$ $0.631$ $158.6$ $0.986$
SCA0.1 $\mathbf{6.19}$ $\mathbf{0.633}$ $\mathbf{287.9}$ $\mathbf{0.984}$
SCA0.25 $\mathbf{5.83}$ $\mathbf{0.607}$ $\mathbf{289.3}$ $\mathbf{0.983}$
SCA0.5 $\mathbf{5.74}$ $\mathbf{0.604}$ $\mathbf{299.6}$ $\mathbf{0.977}$
Fashion No-Defense $8.91$ $0.147$ $235.5$ $0.886$
MNIST Gaussian-Noise $8.67$ $0.132$ $239.8$ $0.815$
GAN $8.66$ $0.147$ $243.3$ $0.883$
Gong et al.[25]++ $8.73$ $0.130$ $220.2$ $0.906$
Titcombe et al.[75] $8.56$ $0.134$ $229.8$ $0.905$
Gong et al.[25] $8.57$ $0.143$ $244.3$ $0.888$
Peng et al.[62] $8.85$ $0.147$ $227.5$ $0.845$
Hayes et al.[28] $8.63$ $0.139$ $218.4$ $0.752$
Wang et al.[82] $8.90$ $0.119$ $210.3$ $0.880$
Sparse-Standard $8.71$ $0.135$ $223.3$ $0.879$
SCA0.1 $\mathbf{8.49}$ $\mathbf{0.039}$ $\mathbf{222.8}$ $\mathbf{0.897}$
SCA0.25 $\mathbf{8.49}$ $\mathbf{0.032}$ $\mathbf{229.9}$ $\mathbf{0.887}$
SCA0.5 $\mathbf{8.45}$ $\mathbf{0.047}$ $\mathbf{233.5}$ $\mathbf{0.876}$
CIFAR10 No-Defense $11.94$ $0.381$ $39.38$ $0.821$
Gaussian-Noise $11.88$ $0.365$ $77.92$ $0.626$
GAN $11.86$ $0.369$ $88.39$ $0.596$
Titcombe et al.[75] $10.89$ $0.346$ $79.19$ $0.792$
Gong et al.[25]++ $11.06$ $0.339$ $78.48$ $0.773$
Gong et al.[25] $11.21$ $0.334$ $92.33$ $0.682$
Peng et al.[62] $11.96$ $0.354$ $120.5$ $0.752$
Hayes et al.[28] $11.12$ $0.342$ $142.1$ $0.626$
Wang et al.[82] $11.02$ $0.346$ $142.6$ $0.756$
Sparse-Standard $10.74$ $0.303$ $137.4$ $0.790$
SCA0.1 $\mathbf{10.59}$ $\mathbf{0.305}$ $\mathbf{144.1}$ $\mathbf{0.787}$
SCA0.25 $\mathbf{10.27}$ $\mathbf{0.279}$ $\mathbf{189.9}$ $\mathbf{0.772}$
SCA0.5 $\mathbf{10.23}$ $\mathbf{0.276}$ $\mathbf{189.7}$ $\mathbf{0.744}$
@@
MNIST No-Defense $40.87$ $0.982$ $16.31$ $0.971$
Gaussian-Noise $40.88$ $0.983$ $15.88$ $0.958$
GAN $40.69$ $0.981$ $16.59$ $0.968$
Titcombe et al.[75] $31.18$ $0.863$ $47.32$ $0.980$
Gong et al.[25]++ $30.37$ $0.838$ $72.99$ $0.987$
Gong et al.[25] $29.05$ $0.817$ $75.39$ $0.985$
Peng et al.[62] $18.44$ $0.354$ $111.6$ $0.968$
Hayes et al.[28] $19.75$ $0.488$ $298.8$ $0.871$
Wang et al.[82] $27.26$ $0.862$ $72.66$ $0.962$
Sparse-Standard $21.34$ $0.439$ $142.9$ $0.986$
SCA0.1 $\mathbf{19.54}$ $\mathbf{0.502}$ $\mathbf{178.5}$ $\mathbf{0.984}$
SCA0.25 $\mathbf{18.81}$ $\mathbf{0.340}$ $\mathbf{174.1}$ $\mathbf{0.983}$
SCA0.5 $\mathbf{17.85}$ $\mathbf{0.164}$ $\mathbf{335.5}$ $\mathbf{0.977}$
Fashion No-Defense $37.86$ $0.975$ $13.91$ $0.886$
MNIST Gaussian-Noise $36.54$ $0.969$ $16.49$ $0.815$
GAN $37.68$ $0.974$ $19.26$ $0.883$
Gong et al.[25]++ $27.71$ $0.794$ $41.35$ $0.906$
Titcombe et al.[75] $26.66$ $0.759$ $53.76$ $0.905$
Gong et al.[25] $21.24$ $0.523$ $93.08$ $0.888$
Peng et al.[62] $17.98$ $0.368$ $70.53$ $0.880$
Hayes et al.[28] $21.13$ $0.297$ $223.3$ $0.752$
Wang et al.[82] $25.98$ $0.806$ $41.87$ $0.838$
Sparse-Standard $19.35$ $0.446$ $128.4$ $0.879$
SCA0.1 $\mathbf{17.92}$ $\mathbf{0.209}$ $\mathbf{196.1}$ $\mathbf{0.897}$
SCA0.25 $\mathbf{17.03}$ $\mathbf{0.186}$ $\mathbf{195.2}$ $\mathbf{0.887}$
SCA0.5 $\mathbf{14.51}$ $\mathbf{0.069}$ $\mathbf{423.2}$ $\mathbf{0.876}$
CIFAR10 No-Defense $21.17$ $0.477$ $70.96$ $0.821$
Gaussian-Noise $20.26$ $0.220$ $77.42$ $0.626$
GAN $19.71$ $0.259$ $132.0$ $0.596$
Titcombe et al.[75] $18.62$ $0.174$ $171.9$ $0.792$
Gong et al.[25]++ $18.27$ $0.209$ $149.1$ $0.773$
Gong et al.[25] $19.10$ $0.150$ $133.8$ $0.682$
Peng et al.[62] $17.20$ $0.002$ $130.3$ $0.717$
Hayes et al.[28] $17.95$ $0.002$ $142.4$ $0.626$
Wang et al.[82] $17.08$ $0.002$ $136.1$ $0.793$
Sparse-Standard $18.01$ $0.003$ $168.6$ $0.790$
SCA0.1 $\mathbf{17.09}$ $\mathbf{0.001}$ $\mathbf{172.0}$ $\mathbf{0.787}$
SCA0.25 $\mathbf{16.78}$ $\mathbf{0.001}$ $\mathbf{189.3}$ $\mathbf{0.772}$
SCA0.5 $\mathbf{16.24}$ $\mathbf{0.001}$ $\mathbf{197.0}$ $\mathbf{0.744}$
@@
MNIST No-Defense $31.21$ $0.923$ $19.64$ $0.963$
Gaussian-Noise $31.07$ $0.922$ $23.27$ $0.972$
GAN $28.39$ $0.894$ $27.26$ $0.969$
Gong et al.[25] $28.30$ $0.806$ $69.38$ $0.986$
Titcombe et al.[75] $25.40$ $0.713$ $76.88$ $0.952$
Gong et al.[25]++ $21.94$ $0.591$ $97.33$ $0.991$
Peng et al.[62] $16.90$ $0.475$ $103.2$ $0.960$
Hayes et al.[28] $17.23$ $0.030$ $288.1$ $0.856$
Wang et al.[82] $21.87$ $0.696$ $53.09$ $0.903$
Sparse-Standard $18.71$ $0.288$ $188.4$ $0.981$
SCA0.1 $\mathbf{16.17}$ $\mathbf{0.109}$ $\mathbf{227.4}$ $\mathbf{0.988}$
SCA0.25 $\mathbf{17.40}$ $\mathbf{0.058}$ $\mathbf{301.6}$ $\mathbf{0.980}$
SCA0.5 $\mathbf{14.98}$ $\mathbf{0.044}$ $\mathbf{307.7}$ $\mathbf{0.975}$
Fashion No-Defense $29.66$ $0.911$ $14.33$ $0.868$
MNIST Gaussian-Noise $29.49$ $0.909$ $14.81$ $0.871$
GAN $26.03$ $0.849$ $19.33$ $0.885$
Gong et al.[25] $23.70$ $0.631$ $97.52$ $0.884$
Titcombe et al.[75] $20.48$ $0.565$ $81.01$ $0.872$
Gong et al.[25]++ $25.77$ $0.726$ $57.72$ $0.908$
Peng et al.[62] $20.67$ $0.583$ $46.48$ $0.865$
Hayes et al.[28] $20.10$ $0.256$ $200.6$ $0.748$
Wang et al.[82] $24.53$ $0.588$ $81.79$ $0.881$
Sparse-Standard $19.54$ $0.405$ $200.5$ $0.882$
SCA0.1 $\mathbf{18.11}$ $\mathbf{0.154}$ $\mathbf{171.1}$ $\mathbf{0.904}$
SCA0.25 $\mathbf{17.74}$ $\mathbf{0.188}$ $\mathbf{203.8}$ $\mathbf{0.896}$
SCA0.5 $\mathbf{17.15}$ $\mathbf{0.134}$ $\mathbf{270.4}$ $\mathbf{0.879}$
CIFAR10 No-Defense $16.48$ $0.709$ $47.77$ $0.823$
Gaussian-Noise $14.79$ $0.311$ $149.5$ $0.598$
GAN $14.87$ $0.296$ $13.01$ $0.675$
Titcombe et al.[75] $14.68$ $0.244$ $157.3$ $0.779$
Gong et al.[25]++ $13.32$ $0.003$ $162.4$ $0.691$
Gong et al.[25] $14.55$ $0.291$ $152.1$ $0.644$
Peng et al.[62] $17.18$ $0.002$ $169.1$ $0.707$
Hayes et al.[28] $15.44$ $0.005$ $204.5$ $0.596$
Wang et al.[82] $14.73$ $0.001$ $176.3$ $0.820$
Sparse-Standard $13.22$ $0.003$ $167.9$ $0.769$
SCA0.1 $\mathbf{13.18}$ $\mathbf{0.002}$ $\mathbf{174.2}$ $\mathbf{0.758}$
SCA0.25 $\mathbf{13.07}$ $\mathbf{0.002}$ $\mathbf{181.2}$ $\mathbf{0.742}$
SCA0.5 $\mathbf{12.88}$ $\mathbf{0.002}$ $\mathbf{375.3}$ $\mathbf{0.739}$
@@
MNIST Titcombe et al.[75]-$1.0$ $6.60$ $0.685$ $280.1$ $0.938$
SCA0.1 $\mathbf{6.19}$ $\mathbf{0.633}$ $\mathbf{287.9}$ $\mathbf{0.984}$
SCA0.25 $\mathbf{5.83}$ $\mathbf{0.607}$ $\mathbf{289.3}$ $\mathbf{0.983}$
SCA0.5 $\mathbf{5.74}$ $\mathbf{0.604}$ $\mathbf{299.6}$ $\mathbf{0.977}$
Fashion Titcombe et al.[75]-$1.0$ $8.72$ $0.1412$ $232.1$ $0.823$
MNIST SCA0.1 $\mathbf{8.49}$ $\mathbf{0.039}$ $\mathbf{222.8}$ $\mathbf{0.897}$
SCA0.25 $\mathbf{8.49}$ $\mathbf{0.032}$ $\mathbf{229.9}$ $\mathbf{0.887}$
SCA0.5 $\mathbf{8.45}$ $\mathbf{0.047}$ $\mathbf{233.5}$ $\mathbf{0.876}$
CIAFR10 Titcombe et al.[75]-$1.0$ $10.75$ $0.335$ $112.7$ $0.779$
SCA0.1 $\mathbf{10.59}$ $\mathbf{0.305}$ $\mathbf{144.1}$ $\mathbf{0.787}$
SCA0.25 $\mathbf{10.27}$ $\mathbf{0.279}$ $\mathbf{189.9}$ $\mathbf{0.772}$
SCA0.5 $\mathbf{10.23}$ $\mathbf{0.276}$ $\mathbf{189.7}$ $\mathbf{0.744}$
@@
MNIST Titcombe et al.[75]-$1.0$ $24.89$ $0.664$ $50.64$ $0.938$
SCA0.1 $\mathbf{19.54}$ $\mathbf{0.502}$ $\mathbf{178.5}$ $\mathbf{0.984}$
SCA0.25 $\mathbf{18.81}$ $\mathbf{0.340}$ $\mathbf{174.1}$ $\mathbf{0.983}$
SCA0.5 $\mathbf{17.85}$ $\mathbf{0.164}$ $\mathbf{335.5}$ $\mathbf{0.977}$
Fashion Titcombe et al.[75]-$1.0$ $20.21$ $0.567$ $80.55$ $0.823$
MNIST SCA0.1 $\mathbf{17.92}$ $\mathbf{0.209}$ $\mathbf{196.1}$ $\mathbf{0.897}$
SCA0.25 $\mathbf{17.03}$ $\mathbf{0.186}$ $\mathbf{195.2}$ $\mathbf{0.887}$
SCA0.5 $\mathbf{14.51}$ $\mathbf{0.069}$ $\mathbf{423.2}$ $\mathbf{0.876}$
CIFAR10 Titcombe et al.[75]-$1.0$ $18.71$ $0.672$ $170.8$ $0.779$
SCA0.1 $\mathbf{17.09}$ $\mathbf{0.001}$ $\mathbf{172.0}$ $\mathbf{0.787}$
SCA0.25 $\mathbf{16.78}$ $\mathbf{0.001}$ $\mathbf{189.3}$ $\mathbf{0.772}$
SCA0.5 $\mathbf{16.24}$ $\mathbf{0.001}$ $\mathbf{197.0}$ $\mathbf{0.744}$